OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x,y). x-2y=0 4x-3y=15

OpenStudy (apoorvk):

Same kinda question - try this one yourself - can you?

OpenStudy (anonymous):

I dont understand how to do it. Ive looked everywhere to try to figure out and it doesnt matter how much theyll tell me how to do it, I will never understand because its confusing. And when I try by myself, I always come up with a wrong answer and I get the answer wrong on my schools computer.

OpenStudy (anonymous):

x - 2y = 0 add 2y to both sides x = 2y Now substitute this into the second equation 4x - 3y = 15 4(2y) - 3y = 15 simplify 8y -3y =15 combine like terms 5y = 15 divide by 5 y = 3 Now plug this back into the original equation x - 2y = 0 x - 2(3) = 0 simplify x - 6 = 0 now add 6 to both sides x = 6 So the answer is (6,3) You can check your answer by substituting this into the 2nd equation. 4x - 3y = 15 4(6) - 3(3) = 15 24 - 9 = 15 15=15